Suppose the Three Stooges have different demands for doughnuts. Moe’s demand is given by Q = 15 – 3P; Larry’s demand is...

  1. Suppose the Three Stooges have different demands for doughnuts. Moe’s demand is given by Q = 15 – 3P; Larry’s demand is given by Q = 9–P; Curly’s demand is given by Q = 6 – 3P.
    1. Derive the market demand curve for doughnuts algebraically.
    2. Graph the market demand curve for doughnuts. Pay special attention to any kinks in the market demand! Label price and quantity at these points.
Answer #1

Moes demand: Larrys demand curlys demand: Q = 15-3P aa-p a = 6-3P Masket demand = Qutatac Q = 15-3P 19-P 76-3p Q = 30-7P M